You are here

Beam delivery update: ISAC target plans

header

02 December 2014

First off: The cyclotron is running normally, delivering ~125 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This is scheduled to continue overnight and though the coming days.

With that out of the way: We’re currently delivering 46K on a discretionary basis to GRIFFIN from the damaged UCx target in the east target station. The target and ionizer heaters have been configured to run current through the unbroken end of the target and the ionizer, allowing for partial heating of the target. Yield measurements of francium and other isotopes this afternoon showed an improvement of roughly an order of magnitude over those measured yesterday with a cold target, but that still leaves them ~2 orders of magnitude below what we would normally expect.

GRIFFIN will continue to take beam until we’re ready to switch to another experiment. We’re still sorting out when then might happen as there are questions about this weekend’s scheduled 213Fr implantations for experiment L112. We are planning to deliver francium beams for trapping in the time that was scheduled for S1218 next week.

On an unrelated note, the OLIS multicharge ion source is still set up to produce 76Se for DRAGON. The experiment that was scheduled isn’t going to happen because of problems with the first electric dipole in DRAGON, but we will deliver beam for measurements of selenium charge-state distributions starting Friday.

-- ACM